OpenStudy (anonymous):

Pyramid A is a square pyramid with a base side length of 12 inches and a height of 8 inches. Pyramid B is a square pyramid with a base side length of 48 inches and a height of 32 inches. How many times bigger is the volume of pyramid B than pyramid A?

OpenStudy (campbell_st):

well find the column for each and compare The volume of a pyramid is \[V = \frac{1}{3} \times Base Area \times Height\] and both pyramids have square bases...

OpenStudy (anonymous):

OK thanks.

OpenStudy (campbell_st):

oops volume not column

OpenStudy (anonymous):

I think i got it thanks for the help :)

OpenStudy (campbell_st):

glad to help... I think the volume will be 64 times larger... at a guess

OpenStudy (anonymous):

Yep i think i got that. It is just a review question so no harm done.